OpenStudy (anonymous):

x + y = 6 x - y = 8 Solve the system of equations. (1, -7) (7, -1) no solution

OpenStudy (anonymous):

substitute, (7,-1)

OpenStudy (anonymous):

add the two equations 2x=14 x=7 substitute x=7 in any of the two equations say 1st one and get the value of y 7+y=6 y=6-7 y=-1

OpenStudy (anonymous):

@nitz, thanks that was much more elaborate. i think i actually learned something:P

OpenStudy (anonymous):

u are welcome